OpenStudy (anonymous):

Write the equation of the line parallel to y = x+ 1 through B (4, -3)

OpenStudy (anonymous):

I am not 100% sure on this question, but i think you will be using a formula. A slope formula, which if i am not wrong is y2-y1/ x2-x1 but im sure someone else will give u it.

OpenStudy (gorv):

parallel line have same slope so slope of parallel line = 1 as of give line pass through 4,-3 so equation is y-(-3)=1*(x-4) y+3=x-4 y=x-4-3 y=x-7

OpenStudy (gorv):

@carolyncreighton